# An Approximation Algorithm for Monotone Submodular Cost Allocation

### v1

@@ Abstract
In this paper, we consider the minimum submodular cost allocation (MSCA) problem. The input of MSCA is $k$ non-negative submodular functions $f_{1},\ldots,f_{k}$ on the ground set $N$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $X_{1},\ldots,X_{k}$ so that $\sum_{i=1}^{k}f_{i}(X_{i})$ is minimized. In this paper, we focus on the case when $f_{1},\ldots,f_{k}$ are monotone (denoted by Mono-MSCA). We provide a natural LP-relaxation for Mono-MSCA, which is equivalent to the convex program relaxation introduced by Chekuri and Ene. We show that the integrality gap of the LP-relaxation is at most $k/2$, which yields a $k/2$-approximation algorithm for Mono-MSCA. We also show that the integrality gap of the LP-relaxation is at least $k/2-\epsilon$ for any constant $\epsilon>0$ when $k$ is fixed.
@@ 1 Introduction
Let $N$ be a finite set of cardinality $n$, and $k\geq 2$ a positive integer. A set function $f:2^{N}\to\mathbb{R}$ is called submodular if $f(S)+f(T)\geq f(S\cup T)+f(S\cap T)$ holds for every $S,T\subseteq N$. In this paper, we focus on the Minimum Submodular Cost Allocation (MSCA) problem, introduced by Chekuri and Ene [5]. The input of MSCA is $k$ non-negative submodular functions $f_{1},\ldots,f_{k}:2^{N}\to\mathbb{R}_{+}$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $X_{1},\ldots,X_{k}$ so that $\sum_{i=1}^{k}f_{i}(X_{i})$ is minimized, i.e.,
$\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$
Throughout this paper, we assume that the input functions $f_{1},\ldots,f_{k}$ satisfy $f_{i}(\emptyset)=0$ for each $i=1,\ldots,k$. Several well-studied problems such as the multiway cut problems in graphs and hypergraphs, the submodular multiway partition problems, and uniform metric labeling can be cast as special cases of MSCA. In the case when all submodular functions are monotone, MSCA is equivalent to the submodular facility location problem considered by Svitkina and Tardos [18]. In this paper, we focus on this setting, i.e., the case where each of $f_{1},\ldots,f_{k}$ is monotone submodular. We denote MSCA with monotone submodular functions $f_{1},\ldots,f_{k}$ by Mono-MSCA. The aim of this paper is to investigate the integrality gap of a natural LP-relaxation for Mono-MSCA when $k$ is small.
@@
To provide an LP-relaxation for Mono-MSCA, we introduce some notation. Let $\mathbf{1}\in\mathbb{Z}^{N}_{>0}$ denote the all-ones vector and $k]:=\{1,2,\ldots,k\}$. For any $S\subseteq N$, let $\chi_{S}\in\{0,1\}^{N}$ denote the characteristic vector of $S$. We consider the following linear programming relaxation for MSCA:
$(\textsc{LP-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
@@
In this LP-relaxation, we interpret $y_{i}(S)=1$ to mean that the set $S$ is assigned to $f_{i}$ in MSCA, whereas $y_{i}(S)=0$ means that $S$ is not assigned to $f_{i}$. We note that if an integer feasible solution $(y_{1},\ldots,y_{k})$ satisfies $y_{i}(S_{1})=y_{i}(S_{2})=1$ for two distinct sets $S_{1},S_{2}\subseteq N$ and $i\in[k]$, then by the submodularity of $f_{i}$ we can update $y_{i}(S_{1}):=y_{i}(S_{2}):=0$ and $y_{i}(S_{1}\cup S_{2}):=1$ without increasing the objective value. Hence, after such uncrossing operations, any integer feasible solution to LP-Rel induces a partition of $N$ for MSCA.
The linear program LP-Rel is equivalent to the convex program relaxation introduced by Chekuri and Ene [5] (see Appendix A for details). The linear program LP-Rel can also be seen as the dual linear program of the unweighted $k$-polymatroid intersection (see [16] and Appendix B for details). Regarding the integrality gap of LP-Rel, Ene and Vondrák [8] showed that if $\mathrm{P}\neq\mathrm{NP}$, then the integrality gap of LP-Rel for general submodular functions cannot be upper-bounded by any factor. Hence, it is natural to focus on special cases such as a monotone case that admit a bounded integrality gap. We refer to LP-Rel restricted to monotone submodular functions $f_{1},\ldots,f_{k}$ as Mono-LP-Rel. Svitkina and Tardos [18] gave a $(1+\ln n)$-approximation algorithm for Mono-MSCA, and Chekuri and Ene [5] showed that the integrality gap of Mono-LP-Rel is at most $(1+\ln n)$. Santiago and Shepherd [15] gave a $k$-approximation algorithm for Mono-MSCA and showed that the integrality gap of Mono-LP-Rel is at most $k$. On the negative side, Svitkina and Tardos [18] showed that the set cover problem with a collection of $k$ sets can be reduced to Mono-MSCA. This implies that if $k$ is a part of input, Mono-MSCA cannot be approximated within a factor of $(1-o(1))\cdot\ln n$ [10], which matches the approximation ratio shown by Svitkina and Tardos [18]. On the other hand, if $k\leq\log n$, then the set cover problem with a collection of $k$ sets can be solved in time polynomial in $n$ by brute-force. Hence, in this case the reduction does not yield any approximation lower bound for Mono-MSCA. Therefore, the approximability of Mono-MSCA and the integrality gap of Mono-LP-Rel for small $k$ are much less understood. In this paper, we aim to obtain tight upper and lower bounds of the integrality gap of Mono-LP-Rel when $k$ is small.
@@
As mentioned previously, the current best upper bound on the integrality gap of Mono-LP-Rel is $\min\{k,1+\ln n\}$. Our first result improves this to $k/2$.
@@
The integrality gap of Mono-LP-Rel is at most $k/2$.
When $k=2$, Mono-LP-Rel coincides with the dual linear program of the unweighted polymatroid intersection. Hence, when $k=2$, Theorem 1.1 corresponds to the (total) dual integrality of the polymatroid intersection. Since the proof of Theorem 1.1 is constructive, it yields a $k/2$-approximation algorithm for Mono-MSCA, which improves the current best approximation ratio $\min\{k,1+\ln n\}$.
@@
There exists a $k/2$-approximation algorithm for Mono-MSCA.
Our next result is a lower bound on the integrality gap of Mono-LP-Rel, which is tight when $k$ is fixed.
@@
If $\log n\geq k$, then the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
We note that if $k$ is fixed, Theorem 1.3 implies that the integrality gap of Mono-LP-Rel is at least $k/2-\epsilon$ for any constant $\epsilon>0$, which matches the upper bound in Theorem 1.1. Another consequence is that if $k\leq\sqrt{\log n}$, then the lower bound in Theorem 1.3 satisfies $\frac{k}{2}-\frac{k(k-1)}{\log n}\geq\frac{k}{2}-1$, which almost matches the upper bound $k/2$ in Theorem 1.1.
Our proof of Theorem 1.1 is constructive as follows. We first obtain an optimal solution $(y_{1},\ldots,y_{k})$ to Mono-LP-Rel whose supports are chains via the ellipsoid method. Then, we take the top $2/k$ of a fraction of the largest sets in the chains. We show that such a collection of sets can be partitioned into groups of $k$ sets, each of which has union $N$. The existence of this partition implies that the integrality gap of Mono-LP-Rel is at most $k/2$. In the proof of Theorem 1.3, we construct $k$ monotone submodular functions such that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
@@
Recently, Bi et al. [3] considered the monotone submodular multiway partition problem (Mono-Sub-MP), which is a special case of Mono-MSCA. They showed that Mono-Sub-MP admits a $4/3$-approximation, and does not admit a $(10/9-\epsilon)$-approximation for every constant $\epsilon>0$. For further recent works related to submodular partition problems, see [1, 2, 4, 11, 14].
@@
The rest of the paper is organized as follows. In Section 2, we show that the integrality gap of Mono-LP-Rel is at most $k/2$, which yields a $k/2$-approximation algorithm for Mono-MSCA. In Section 3, we show that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$.
@@ 2 Approximation algorithm
In this section, we show that the integrality gap of Mono-LP-Rel is at most $k/2$, which yields a $k/2$-approximation algorithm for Mono-MSCA.
Let $N$ be a finite set of cardinality $n$, and $k\geq 2$ a positive integer. For each $i\in[k]$, let $f_{i}:2^{N}\to\mathbb{R}_{+}$ be a monotone submodular function with $f_{i}(\emptyset)=0$. Recall that Mono-LP-Rel is defined as follows:
$(\textsc{Mono-LP-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
@@
To prove the upper bound of the integrality gap, we show that an integer solution to Mono-LP-Rel achieving a $k/2$-approximation can be constructed from an optimal solution to Mono-LP-Rel. Since the coefficient matrix of the constraints in Mono-LP-Rel is rational, Mono-LP-Rel has a rational optimal solution $(y^{*}_{1},\ldots,y^{*}_{k})$. Let $\mathcal{L}_{i}:=\{S\subseteq N\mid y^{*}_{i}(S)>0\}$ and $d_{i}:=|\mathcal{L}_{i}|$ for each $i\in[k]$. By the uncrossing operations, we may assume that $\mathcal{L}_{i}$ is a chain for each $i\in[k]$. Let $C_{i}^{1}\supsetneq C_{i}^{2}\supsetneq\cdots\supsetneq C_{i}^{d_{i}}$ be the members of $\mathcal{L}_{i}$ for each $i\in[k]$. Let $M$ be the product of k(k1)k(k-1) and the denominators of $y^{*}_{i}(C_{i}^{j})$ for all $i,j$. Then, $M\cdot y^{*}_{i}$ is an integer vector for every $i\in[k]$. For each $i\in[k]$ and $j\in[M]$, define $U_{i}^{j}\subseteq N$ as follows:
$U_{i}^{j}:=\begin{cases}C_{i}^{\ell}&\mathrm{if\ }1+\displaystyle\sum_{p=1}^{\ell-1}M\cdot y_{i}^{*}(C_{i}^{p})\leq j\leq\displaystyle\sum_{p=1}^{\ell}M\cdot y_{i}^{*}(C_{i}^{p})\ \mathrm{for\ some\ }\ell\in[d_{i}],\\ \emptyset&\mathrm{if\ }1+\displaystyle\sum_{p=1}^{d_{i}}M\cdot y_{i}^{*}(C_{i}^{p})\leq j\leq M.\end{cases}$
Intuitively, for each $i\in[k]$, we consider the multiset that contains $M\cdot y^{*}_{i}(C_{i}^{\ell})$ copies of $C_{i}^{\ell}$ for all $\ell\in[d_{i}]$, and $U_{i}^{j}$ is the $j$-th largest set in the multiset for each $j\in[M]$. By the definition of $U_{i}^{j}$, we have the following lemma.
@@
For each $i\in[k]$ and $\ell=1,\ldots,d_{i}$, the number of $j\in[M]$ with $U_{i}^{j}=C_{i}^{\ell}$ is $M\cdot y_{i}^{*}(C_{i}^{\ell})$.
We now aim to construct an integer solution to Mono-LP-Rel (i.e., a partition of $N$) from $U_{i}^{j}$s that achieves $k/2$-approximation. To this end, we prepare the following two lemmas.
@@
By Lemma 2.1, we have
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)=\sum_{S\in\mathcal{L}_{i},i\in[k]}y^{*}_{i}(S)f_{i}(S)=\sum_{1\leq\ell\leq d_{i},i\in[k]}y^{*}_{i}(C_{i}^{\ell})f_{i}(C_{i}^{\ell})$
$=\sum_{1\leq\ell\leq d_{i},i\in[k]}\dfrac{1}{M}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|\cdot f_{i}(C_{i}^{\ell})$
$=\sum_{\begin{subarray}{c}i\in[k],j\in[M]\\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }1\leq\ell\leq d_{i}\end{subarray}}\dfrac{1}{M}f_{i}(U_{i}^{j})=\sum_{i\in[k],j\in[M]}\dfrac{1}{M}f_{i}(U_{i}^{j}).$
The last equality follows since if $U_{i}^{j}\neq C_{i}^{\ell}$ for all $\ell=1,\ldots,d_{i}$, then we have $U_{i}^{j}=\emptyset$ by the definition. ∎
@@
Let $a_{1},\ldots,a_{k}\in[M]$ be positive integers such that $\sum_{i=1}^{k}a_{i}\leq M+k-1$. Then, $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$.
@@
Suppose to the contrary that $\bigcup_{i=1}^{k}U_{i}^{a_{i}}\neq N$. Then, there exists $e\in N$ such that $e\notin U_{i}^{a_{i}}$ for every $i\in[k]$. Combined with $U_{i}^{1}\supseteq U_{i}^{2}\supseteq\cdots\supseteq U_{i}^{M}$ for each $i\in[k]$, this implies $e\notin U_{i}^{j}$ for every $i\in[k]$ and integer $j$ with $a_{i}\leq j\leq M$. Hence, by Lemma 2.1 we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}M\cdot y^{*}_{i}(S)=\sum_{\begin{subarray}{c}S\in\mathcal{L}_{i},e\in S\\ i\in[k]\end{subarray}}M\cdot y^{*}_{i}(S)=\sum_{\begin{subarray}{c}1\leq\ell\leq d_{i},e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}M\cdot y^{*}_{i}(C_{i}^{\ell})$
$=\sum_{\begin{subarray}{c}1\leq\ell\leq d_{i},e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|$
$=\sum_{i\in[k]}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}],\ e\in U_{i}^{j}\right\}\right|$
$\leq\sum_{i\in[k]}\left|\left\{j\in[a_{i}-1]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}]\right\}\right|$
@@
Since $y^{*}_{1},\ldots,y^{*}_{k}$ is a feasible solution to Mono-LP-Rel, we have
@@
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}M\cdot y_{i}^{*}(S)=M,$

### v2

@@ Abstract
In this paper, we consider the minimum submodular cost allocation (MSCA) problem. The input of MSCA is $k$ non-negative submodular functions $f_{1},f_{2},\ldots,f_{k}$ on the ground set $N$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $S_{1},S_{2},\ldots,S_{k}$ so that $\sum_{i=1}^{k}f_{i}(S_{i})$ is minimized. In this paper, we focus on the case when $f_{1},f_{2},\ldots,f_{k}$ are monotone, which coincides with the submodular facility location problem considered by Svitkina and Tardos. We show that the integrality gap of a natural LP-relaxation for MSCA with monotone submodular functions is at most $k/2$, yielding a $k/2$-approximation algorithm. We also prove a nearly matching lower bound: the integrality gap is at least $k/2-\epsilon$ for any constant $\epsilon>0$ when $k$ is fixed.
@@ 1 Introduction
Let $N$ be a finite set and $k\geq 2$ a positive integer. A set function $f:2^{N}\to\mathbb{R}$ is called submodular if $f(S)+f(T)\geq f(S\cup T)+f(S\cap T)$ holds for every $S,T\subseteq N$. In this paper, we focus on the Minimum Submodular Cost Allocation (MSCA) problem, introduced by Chekuri and Ene [5]. The input of MSCA is $k$ non-negative submodular functions $f_{1},f_{2},\ldots,f_{k}:2^{N}\to\mathbb{R}_{+}$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $S_{1},S_{2},\ldots,S_{k}$ so that $\sum_{i=1}^{k}f_{i}(S_{i})$ is minimized, i.e.,
$\min\left\{\sum_{i=1}^{k}f_{i}(S_{i})\ \middle|\ S_{1},S_{2},\ldots,S_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$
Several well-studied problems such as the multiway cut problems in graphs and hypergraphs, the submodular multiway partition problems, and uniform metric labeling can be cast as special cases of MSCA [5]. When all submodular functions are monotone, MSCA is equivalent to the submodular facility location problem considered by Svitkina and Tardos [18]. In this paper, we focus on this setting, that is, we assume that $f_{i}(S)\leq f_{i}(T)$ for all $S\subseteq T\subseteq N$ and $i=1,2,\ldots,k$. We call this monotone case of MSCA the Minimum Monotone Submodular Cost Allocation (MMSCA) problem. The aim of this paper is to investigate the integrality gap of a natural LP-relaxation for MMSCA when $k$ is small.
@@
To provide an LP-relaxation for MMSCA, we introduce some notation. Let $\mathbf{1}\in\mathbb{Z}^{N}_{>0}$ denote the all-ones vector and $k]:=\{1,2,\ldots,k\}$. For any $S\subseteq N$, let $\chi_{S}\in\{0,1\}^{N}$ denote the characteristic vector of $S$. Throughout this paper, we assume that the input functions $f_{1},f_{2},\ldots,f_{k}$ satisfy $f_{i}(\emptyset)=0$ for each $i=1,2,\ldots,k$. We consider the following linear programming relaxation for MSCA:
$(\textsc{LP})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
@@
In this LP-relaxation, we interpret $y_{i}(S)=1$ to mean that $S$ is assigned to $f_{i}$ in MSCA, whereas $y_{i}(S)=0$ means that $S$ is not assigned to $f_{i}$. We note that if an integer feasible solution $(y_{1},y_{2},\ldots,y_{k})$ satisfies $y_{i}(S)=y_{i}(T)=1$ for two distinct sets $S,T\subseteq N$ and $i\in[k]$, then by the submodularity of $f_{i}$ we can update $y_{i}(S):=y_{i}(T):=0$ and $y_{i}(S\cup T):=1$ without increasing the objective value. Hence, after such uncrossing operations, any integer feasible solution to (LP) induces a partition of $N$ for MSCA.
The linear program (LP) is equivalent to the convex program relaxation introduced by Chekuri and Ene [5]. The linear program (LP) can also be seen as the dual linear program of the unweighted $k$-polymatroid intersection; see, e.g., [16]. Regarding the integrality gap of (LP), Ene and Vondrák [8] showed that if $\mathrm{P}\neq\mathrm{NP}$, then the integrality gap of (LP) for general submodular functions cannot be upper-bounded by any factor. Hence, it is natural to focus on special cases such as a monotone case that admit a bounded integrality gap. We refer to (LP) restricted to monotone submodular functions $f_{1},f_{2},\ldots,f_{k}$ as $\mathrm{(LP_{mono})}$. Svitkina and Tardos [18] gave a $(1+\ln n)$-approximation algorithm for MMSCA, and Chekuri and Ene [5] showed that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $(1+\ln n)$, where $n:=|N|$. Santiago and Shepherd [15] showed that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k$, yielding a $k$-approximation for MMSCA. On the negative side, Svitkina and Tardos [18] showed that the set cover problem with a collection of $k$ sets can be reduced to MMSCA. This implies that if $k$ is a part of input, MMSCA cannot be approximated within a factor of $(1-o(1))\cdot\ln n$ [10], which matches the approximation ratio shown by Svitkina and Tardos [18]. On the other hand, if $k\leq\log n$, then the set cover problem with a collection of $k$ sets can be solved in time polynomial in $n$ by brute-force. Hence, in this case the reduction does not yield any approximation lower bound for MMSCA. Therefore, the approximability of MMSCA and the integrality gap of $\mathrm{(LP_{mono})}$ for small $k$ are much less understood. In this paper, we aim to obtain tight upper and lower bounds of the integrality gap of $\mathrm{(LP_{mono})}$ when $k$ is small.
@@
As mentioned previously, the current best upper bound on the integrality gap of $\mathrm{(LP_{mono})}$ is $\min\{k,1+\ln n\}$. Our first result improves this to $k/2$.
@@
The integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$.
Recall that when $k=2$, $\mathrm{(LP_{mono})}$ is the dual linear program of the unweighted polymatroid intersection. Since the constraint system of the unweighted polymatroid intersection is totally dual integral (see, e.g., [16]), its dual admits an integral optimal solution, which matches Theorem 1.1 for $k=2$. Hence, Theorem 1.1 can be viewed as extending the (total) dual integrality of polymatroid intersection, in the sense that it yields a $k/2$ bound on the dual integrality gap for $k$-polymatroid intersection.
Since the proof of Theorem 1.1 is constructive, it yields a $k/2$-approximation algorithm for MMSCA, which improves the current best approximation ratio $\min\{k,1+\ln n\}$.
@@
There exists a $k/2$-approximation algorithm for MMSCA.
Our next result is a lower bound on the integrality gap of $\mathrm{(LP_{mono})}$, which is tight when $k$ is fixed.
@@
If $\log n\geq k$, then the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
We note that if $k$ is fixed, Theorem 1.3 implies that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $k/2-\epsilon$ for any constant $\epsilon>0$, which matches the upper bound in Theorem 1.1. Another consequence is that if $k\leq\sqrt{\log n}$, then the lower bound in Theorem 1.3 satisfies $\frac{k}{2}-\frac{k(k-1)}{\log n}\geq\frac{k}{2}-1$, which almost matches the upper bound $k/2$ in Theorem 1.1.
Our proof of Theorem 1.1 is constructive as follows. We first obtain an optimal solution $(y_{1},y_{2},\ldots,y_{k})$ to $\mathrm{(LP_{mono})}$ whose supports are chains via the ellipsoid method. Then, we take the top $2/k$ of a fraction of the largest sets in the chains. We show that such a collection of sets can be partitioned into groups of $k$ sets, each of which has union $N$. The existence of this partition implies that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$. In the proof of Theorem 1.3, we construct $k$ monotone submodular functions such that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
@@
Svitkina and Tardos [18] considered the submodular facility location problem defined as follows. The input consists of a set of facilities $\mathcal{F}$, a set of demands or clients $\mathcal{D}$, a non-negative cost $c_{ij}$ to connect facility $i$ to client $j$ for each $i\in\mathcal{F}$ and $j\in\mathcal{D}$, and a facility opening cost function $f_{i}$ for each $i\in\mathcal{F}$. The goal is to assign each client to a facility so as to minimize the sum of the facility opening cost and the connection cost, where the facility opening cost for a facility $i$ is a monotone submodular function $f_{i}$ of the clients assigned to $i$. Since the sum of the facility opening cost and the connection cost for a facility $i$ is a monotone submodular function of the clients assigned to $i$, the submodular facility location problem is equivalent to MMSCA. As mentioned above, Svitkina and Tardos [18] gave a $(1+\ln|\mathcal{D}|)$-approximation algorithm for the submodular facility location problem, and showed a matching lower bound $(1-o(1))\cdot\ln|\mathcal{D}|$ when $|\mathcal{F}|$ is not fixed or bounded, while our result improves the approximation ratio to $|\mathcal{F}|/2$.
Recently, Bi et al. [3] considered the submodular multiway partition problem for monotone submodular functions, which they call monotone submodular multiway partition (Mono-Sub-MP). In Mono-Sub-MP, we are given a non-negative monotone submodular function $f:2^{N}\to\mathbb{R}_{+}$ with $f(\emptyset)=0$ and $k$ terminals $t_{1},t_{2},\ldots,t_{k}\in N$. The goal is to find a partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ with $t_{i}\in S_{i}$ for every $i\in[k]$ in order to minimize $\sum_{i=1}^{k}f(S_{i})$. As the submodular multiway partition problem can be seen as a special case of MSCA, Mono-Sub-MP is a special case of MMSCA as follows. Define $N^{\prime}:=N\setminus\{t_{1},t_{2},\ldots,t_{k}\}$ and $f_{i}(S):=f(S\cup\{t_{i}\})-f(\{t_{i}\})$ for each $i\in[k]$ and $S\subseteq N^{\prime}$. Then, $f_{i}$ is a non-negative monotone submodular functions with $f_{i}(\emptyset)=0$ for every $i\in[k]$. Hence, Mono-Sub-MP can be cast as a special case of MMSCA. Bi et al. [3] showed that Mono-Sub-MP admits a $4/3$-approximation, and does not admit a $(10/9-\epsilon)$-approximation for every constant $\epsilon>0$. For further recent works related to submodular partition problems, see [1, 2, 4, 11, 14].
@@
The rest of the paper is organized as follows. In Section 2, we show that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$, which yields a $k/2$-approximation algorithm for MMSCA. In Section 3, we show that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$.
@@ 2 Approximation algorithm
In this section, we prove Theorems 1.1 and 1.2, showing that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$, which yields a $k/2$-approximation algorithm for MMSCA.
Let $N$ be a finite set of cardinality $n$, and $k\geq 2$ a positive integer. For each $i\in[k]$, let $f_{i}:2^{N}\to\mathbb{R}_{+}$ be a monotone submodular function with $f_{i}(\emptyset)=0$. Recall that $\mathrm{(LP_{mono})}$ is defined as follows:
$\textsc{$\mathrm{(LP_{mono})}$}\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
@@
To prove the upper bound of the integrality gap, we show that a $k/2$-approximate integer solution to $\mathrm{(LP_{mono})}$ can be constructed from an optimal solution to $\mathrm{(LP_{mono})}$. Since the coefficient matrix of the constraints in $\mathrm{(LP_{mono})}$ is rational, $\mathrm{(LP_{mono})}$ has a rational optimal solution $(y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k})$. Let $\mathcal{L}_{i}:=\{S\subseteq N\mid y^{*}_{i}(S)>0\}$ and $d_{i}:=|\mathcal{L}_{i}|$ for each $i\in[k]$. By the uncrossing operations, we may assume that $\mathcal{L}_{i}$ is a chain for each $i\in[k]$. Let $C_{i}^{1}\supsetneq C_{i}^{2}\supsetneq\cdots\supsetneq C_{i}^{d_{i}}$ be the members of $\mathcal{L}_{i}$ for each $i\in[k]$.
For convenience of discussion, we will convert $y^{*}_{i}$ into an integer vector $z_{i}$ by multiplying it by some constant factor, and make $z_{i}(C_{i}^{\ell})$ copies of $C_{i}^{\ell}$ for each $\ell$. To this end, we define $M$ as the product of k(k1)k(k-1) and the denominators of $y^{*}_{i}(C_{i}^{j})$ for all $i,j$. Then, $z_{i}:=M\cdot y^{*}_{i}$ is an integer vector for every $i\in[k]$. For each $i\in[k]$, we define $U_{i}^{1},U_{i}^{2},\ldots,U_{i}^{M}\subseteq N$ by setting $U_{i}^{j}:=C_{i}^{1}$ for the first $z_{i}(C_{i}^{1})$ indices $j=1,2,\ldots,z_{i}(C_{i}^{1})$, and then setting $U_{i}^{j}:=C_{i}^{2}$ for the next $z_{i}(C_{i}^{2})$ indices $j=z_{i}(C_{i}^{1})+1,z_{i}(C_{i}^{1})+2,\ldots,z_{i}(C_{i}^{1})+z_{i}(C_{i}^{2})$, and continue in this way up to $C_{i}^{d_{i}}$. For the remaining indices $j$, setting $U_{i}^{j}:=\emptyset$. In other words, for each $\ell\in[d_{i}]$, we set $U_{i}^{j}:=C_{i}^{\ell}$ for all $j$ in the following block of indices:
$1+\displaystyle\sum_{p=1}^{\ell-1}z_{i}(C_{i}^{p})\leq j\leq\displaystyle\sum_{p=1}^{\ell}z_{i}(C_{i}^{p}).$
By the definition of $U_{i}^{j}$, we have the following lemma.
@@
For each $i\in[k]$ and $\ell\in[d_{i}]$, the number of $j\in[M]$ such that $U_{i}^{j}=C_{i}^{\ell}$ is $z_{i}(C_{i}^{\ell})$.
To show that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$, we will construct a partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ that corresponds to a $k/2$-approximate solution to $\mathrm{(LP_{mono})}$. Such a partition can be constructed by taking the top $2/k$ fraction of the largest sets among $U_{i}^{j}$ for each $i\in[k]$, and partitioning them into groups of $k$ sets whose union is $N$. For the construction, we prepare the following two lemmas.
@@
Since $z_{i}(C_{i}^{\ell})$ is the number of $j\in[M]$ with $U_{i}^{j}=C_{i}^{\ell}$ by Lemma 2.1, we have
$\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S)=\sum_{S\in\mathcal{L}_{i},i\in[k]}z_{i}(S)f_{i}(S)=\sum_{\ell\in[d_{i}],i\in[k]}z_{i}(C_{i}^{\ell})f_{i}(C_{i}^{\ell})$
$=\sum_{\ell\in[d_{i}],i\in[k]}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|\cdot f_{i}(C_{i}^{\ell})$
$=\sum_{\begin{subarray}{c}i\in[k],j\in[M]\\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}]\end{subarray}}f_{i}(U_{i}^{j})=\sum_{i\in[k],j\in[M]}f_{i}(U_{i}^{j}).$
The last equality follows since if $U_{i}^{j}\neq C_{i}^{\ell}$ for all $\ell=1,2,\ldots,d_{i}$, then we have $U_{i}^{j}=\emptyset$ by the definition. ∎
@@
Let $a_{1},a_{2},\ldots,a_{k}\in[M]$ be positive integers such that $\sum_{i=1}^{k}a_{i}\leq M+k-1$. Then, $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$.
@@
Suppose to the contrary that $\bigcup_{i=1}^{k}U_{i}^{a_{i}}\neq N$. Then, there exists $e\in N$ such that $e\notin U_{i}^{a_{i}}$ for every $i\in[k]$. Combined with $U_{i}^{1}\supseteq U_{i}^{2}\supseteq\cdots\supseteq U_{i}^{M}$ for each $i\in[k]$, this implies that $e\notin U_{i}^{j}$ for every $i\in[k]$ and integer $j$ with $a_{i}\leq j\leq M$. Hence, since $z_{i}(C_{i}^{\ell})$ is the number of $j\in[M]$ with $U_{i}^{j}=C_{i}^{\ell}$ by Lemma 2.1, we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=\sum_{\begin{subarray}{c}S\in\mathcal{L}_{i},e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=\sum_{\begin{subarray}{c}\ell\in[d_{i}],e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}z_{i}(C_{i}^{\ell})$
$=\sum_{\begin{subarray}{c}\ell\in[d_{i}],e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|$
$=\sum_{i\in[k]}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}]\ \mathrm{with\ }e\in C_{i}^{\ell}\right\}\right|$
$=\sum_{i\in[k]}\left|\left\{j\in[a_{i}-1]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}]\ \mathrm{with\ }e\in C_{i}^{\ell}\right\}\right|$
@@
Since $(y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k})$ is a feasible solution to $\mathrm{(LP_{mono})}$, we have
@@
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=M,$